Keil is going to make 13 pounds of mixed nuts for a party. Peanuts cost $3.50 per pound and fancy nuts cost $7.50 per pound. If Keil can spend $85.50 on nuts, how many pounds of each should he buy?
Answer by mr.barrett(5)

The question is asking "how many pounds of each should she buy?" That tells us what our variables should be.


x=pounds of peanuts
y=pounds of fancy nuts


x + y = 13 (equation 1)
3.50x + 7.50y = 85.50 (equation 2)


Solve this system of equations. I'll do it by substitution here.
From equation 1, y = 13 - x.
By substitution into equation 2, 3.50x + 7.50(13 - x) = 85.50


Solving for x we get x=3.


Substitute 3 for x into equation 1 to solve for y, and y=13.